# PriorDiffusion: Leverage Language Prior in Diffusion Models for Monocular Depth Estimation

### v2

@@ 1 Introduction
Shown in Figure 1, an emerging solution lies in leveraging language as a complementary modality to provide a prior to resolve ambiguity and visual nuisance. Language descriptions provided by humans offer a wealth of priors about spatial relationships, scene types, object sizes, and depth hierarchies.
In cases where an object is small, partially occluded, visually similar to the background, or its type is ambiguous from a single image, language can indicate the object’s identity and semantics, enabling the model to infer its typical 3D size and shape and make more precise depth predictions for the observed portion. Language inputs are also valuable in resolving texture ambiguities. Descriptive cues can specify the type and function of surfaces in a scene, aiding the model in inferring geometric properties. For instance, if a uniformly textured surface is described as a “white ceiling”, the model can interpret it as a continuous, flat plane at a consistent depth above the observer, extending into the distance. Similarly, for surfaces with repetitive patterns, such as a red and black interwoven carpet, language inputs like “A red and black interwoven carpet extending across the room” can help the model recognize that the repeating pattern belongs to a flat surface, preventing misinterpretation of edges between tiles as depth discontinuities. As shown in Figure 2, better perception can be achieved for insignificant regions indicated in the language description.
@@ 2 Related Work
Monocular depth estimation. Monocular depth estimation is a dense regression problem, requiring pixel-wise predictions of depth from a single image, capturing complex spatial relationships within the scene. Some depth models learn to infer pixel-wise depth in metric scale (i.e. meters) [2, 18, 39, 48, 81, 97, 106, 98] by minimizing loss between depth predictions and ground-truth depth maps. Each model typically applies to only one data domain in which it is trained, with similar camera parameters and object scales.
When ground-truth depth is not available, self-supervised approaches [3, 17, 36, 40, 41, 53, 57, 75, 77, 78, 80, 85, 96, 102, 107, 108, 112, 115] rely on geometric constraints, where scale is attributed through radar [73], inertials [30, 79], lidar [16, 51, 56, 84, 83, 82, 86, 93], or binocular images [24, 28, 27]. To enable generalization across diverse domains, empowering depth estimation in the wild, other works have explored affine-invariant depth estimation [100, 90, 63, 62, 38, 91, 95, 13, 94, 47], which predicts depth up to an unknown global shift and scale. This approach can accommodate varied scenes with different scales, while still preserving the geometric relationships between different objects or regions within the scene. HND [100] hierarchically normalizes the depth representations with spatial information and depth distributions. Depth Anything  [90] learns from large-scale automatically annotated data. DPT [63] leverages vision transformers using a scale- and shift-invariant trimmed loss. MiDas [62] mixes multiple datasets with training objectives invariant to depth range and scale. Depth Anything [90] and Depth Anything v2 [91] train a depth foundation model using large-scale unlabeled data. Existing methods try to predict depth solely from images, which suffers from ambiguity like scale, object orientation, occlusion, and visual nuisance like viewpoints, illumination, appearance, texture, etc. On the other hand, our PriorDiffusion leverages language prior in Diffusion Models to provide a thorough understanding of the scene associated with the input text, to resolve the ambiguity and the nuisance variables in the visual algorithm.
@@
Language for monocular depth estimation. Vision-Language models  [5, 46, 45, 55, 60] acquire a comprehensive understanding of languages and images through pre-training under diverse datasets, thereby establishing a robust foundation for downstream tasks. [88, 89, 104, 7, 105, 106, 117, 98, 99, 11]. Typically, CLIP [60] employs contrastive learning on text-image pairs, enabling a range of applications like few-shot image classification  [21, 103, 59, 114], image segmentation  [64, 111], object detection  [64, 116], and 3D perception [35, 105, 106, 117, 113, 101, 8, 32]. In light of their emerging ability, some works [1, 35, 106, 109] have tried to apply vision-language models for monocular depth estimation. Particularly, DepthCLIP [106] leverages the semantic depth response of CLIP with a depth projection scheme to conduct zero-shot monocular depth estimation. WorDepth [98] learns the variational prior distribution of 3D scenes from text. RSA [99] predicts scale using language description to align relative depth to metric scale. While language has been applied in various aspects of 3D perception, our work is the first to use language prior for diffusion models to enhance 3D perception.
@@ 3 Method
Problem setup. Given an RGB image $x:\Omega\subset\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$, the goal of monocular depth estimation is to predict a dense depth map $y:\Omega\subset\mathbb{R}^{2}\rightarrow\mathbb{R}_{+}$ using a parameterized function $h$, represented as a neural network, such that $y:=h(\cdot)$. We utilize a supervised dataset $\mathcal{D}=\{x^{(m)},\textbf{c}^{(m)},y^{*(m)}\}_{m=1}^{M}$ containing $M$ samples, where $y^{*}:\Omega\subset\mathbb{R}^{2}\rightarrow\mathbb{R}_{+}$ is the ground-truth depth map, and c represents the text caption that describes the image.
Diffusion formulation. The pipeline of PriorDiffusion is shown in Figure 3. To use a diffusion model for monocular depth estimation, following [12, 37, 69, 70, 110, 38], we adopt the formulation of Denoising Diffusion Probabilistic Model (DDPM) [33]. We define a latent variable $\mathbf{z}_{t}$ $\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$that represents a noisy version of the depth map in the latent space at diffusion timestep $t$. The ground truch depth map $y^{*}$ is encoded into a latent representation $\mathbf{z_{y}}$ as supervision using a frozen VAE encoder $E$: $\mathbf{z_{y}}=E(y^{*})$. The diffusion latent after the whole denoising process $\mathbf{z_{0}}$ is decoded back to the image space using the frozen VAE decoder $D$: $\hat{y}=D(\mathbf{z}_{0})$.
@@ 4 Experiments
Training datasets. We adopt the training scheme of Marigold [38], and train our PriorDiffusion on two synthetic datasets that cover both indoor and outdoor scenes. The first dataset, HyperSim [65], is a photorealistic collection of 461 indoor scenes. From this, we use the official split to select around 54K samples from 365 scenes, excluding any incomplete samples. RGB images and depth maps are resized to a resolution of 480 × 640 pixels, and the original distance measurements, relative to the focal point, are converted to conventional depth values relative to the focal plane. The second dataset, Virtual KITTI [4, 20], includes synthetic street scenes from 5 settings with varied conditions, such as different weather and camera perspectives. We use the Virtual KITTI 2 [4] version. For training, we select four scenes, totaling approximately 20K samples, and crop the images to align with the KITTI benchmark resolution [25], setting the maximum depth to 80 meters.
@@
Implementation details. We implemented our method using PyTorch, employing Stable Diffusion v2 [66] as the backbone and adhering to the original pre-training setup with the $v$-objective [68]. During training, we used the DDPM noise scheduler [33] with 1000 diffusion steps, while at inference time, the DDIM scheduler [74] was applied with 50 sampling steps. Our training process spanned 30000 iterations with a batch size of 32, achieved through gradient accumulation over 16 steps (batch size of 2 per step) to fit a single GPU. We optimized the model using the Adam optimizer with a learning rate of $3\cdot 10^{-5}$ and included random horizontal flipping as data augmentation.
@@ C.2. Ablation for prompts to generate text
This prompt is designed to elicit responses that include essential details, such as the positioning of objects, their interactions, and notable features that may impact depth estimation. Note that all the training data we use is synthetic data, and by emphasizing “assuming it’s a real-world image,” we ensure that the descriptions align with the types of inputs and scenarios the model will encounter.

### v3

@@ 1 Introduction
Shown in Figure 1, an emerging solution lies in leveraging language as a complementary modality to provide a prior to resolve ambiguity and visual nuisance. Language descriptions provided by humans offer a wealth of priors about spatial relationships, scene types, object sizes, and depth hierarchies. In cases where an object is small, partially occluded, visually similar to the background, or its type is ambiguous from a single image, language can indicate the object’s identity and semantics, enabling the model to infer its typical 3D size and shape and make more precise depth predictions for the observed portion. Language inputs are also valuable in resolving texture ambiguities. Descriptive cues can specify the type and function of surfaces in a scene, aiding the model in inferring geometric properties. For instance, if a uniformly textured surface is described as a “white ceiling”, the model can interpret it as a continuous, flat plane at a consistent depth above the observer, extending into the distance. Similarly, for surfaces with repetitive patterns, such as a red and black interwoven carpet, language inputs like “A red and black interwoven carpet extending across the room” can help the model recognize that the repeating pattern belongs to a flat surface, preventing misinterpretation of edges between tiles as depth discontinuities. As shown in Figure 2, better perception can be achieved for insignificant regions indicated in the language description.
@@ 2 Related Work
Monocular depth estimation. Monocular depth estimation is a dense regression problem, requiring pixel-wise predictions of depth from a single image, capturing complex spatial relationships within the scene. Some depth models learn to infer pixel-wise depth in metric scale (i.e. meters) [2, 16, 32, 39, 64, 73, 81, 74] by minimizing loss between depth predictions and ground-truth depth maps. Each model typically applies to only one data domain in which it is trained, with similar camera parameters and object scales. To enable generalization across diverse domains, empowering depth estimation in the wild, other works have explored affine-invariant depth estimation [76, 68, 51, 50, 30, 69, 72, 12, 71, 38], which predicts depth up to an unknown global shift and scale. This approach can accommodate varied scenes with different scales, while still preserving the geometric relationships between different objects or regions within the scene. HND [76] hierarchically normalizes the depth representations with spatial information and depth distributions. Depth Anything  [68] learns from large-scale automatically annotated data. DPT [51] leverages vision transformers using a scale- and shift-invariant trimmed loss. MiDas [50] mixes multiple datasets with training objectives invariant to depth range and scale. Depth Anything [68] and Depth Anything v2 [69] train a depth foundation model using large-scale unlabeled data. Existing methods try to predict depth solely from images, which suffers from ambiguity like scale, object orientation, occlusion, and visual nuisance like viewpoints, illumination, appearance, texture, etc. On the other hand, our PriorDiffusion leverages language prior in Diffusion Models to provide a thorough understanding of the scene associated with the input text, to resolve the ambiguity and the nuisance variables in the visual algorithm.
@@
Language for monocular depth estimation. Vision-Language models  [4, 37, 36, 45, 48] acquire a comprehensive understanding of languages and images through pre-training under diverse datasets, thereby establishing a robust foundation for downstream tasks. [66, 67, 79, 6, 80, 81, 87, 74, 75, 43, 62, 31]. For example, CLIP [48] use contrastive learning on text-image pairs, which allows for various applications. These include few-shot image classification  [19, 78, 47], object detection  [52, 86], image segmentation  [52, 84], and 3D perception [28, 80, 81, 87, 85, 77, 7]. Given their growing capabilities, some research [1, 28, 81, 82] has focused on applying vision-language models to monocular depth estimation. Specifically, DepthCLIP [81] utilizes CLIP’s semantic depth response as an ordinal relationship alongside a depth projection method to conduct zero-shot monocular depth estimation. WorDepth [74] learns the variational prior distribution of 3D scenes from text. RSA [75] predicts scale using language description to align relative depth to metric scale. While language has been applied in various aspects of 3D perception, our work is the first to use language prior for diffusion models to enhance 3D perception.
@@ 3 Method
Problem formulation. Formally, given an RGB image $x:\Omega\subset\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$, the objective of monocular depth estimation is to predict a dense depth map $y:\Omega\subset\mathbb{R}^{2}\rightarrow\mathbb{R}_{+}$ using a parameterized function $h$, typically represented as a neural network, such that $y:=h(x)$. We employ a supervised dataset $\mathcal{D}=\{x^{(m)},\textbf{c}^{(m)},y^{*(m)}\}_{m=1}^{M}$ comprised of $M$ samples, where $y^{*}:\Omega\subset\mathbb{R}^{2}\rightarrow\mathbb{R}_{+}$ denotes the ground-truth depth map, and c signifies the corresponding text caption of the image.
Diffusion formulation. The pipeline of PriorDiffusion is shown in Figure 3. To use a diffusion model for monocular depth estimation, following the formulation of  [11, 29, 57, 58, 83, 30], we adopt the formulation of Denoising Diffusion Probabilistic Model (DDPM) [26] for the forward diffusion process, reverse diffusion process, training objective and inference mentioned below. A latent variable $\mathbf{z}_{t}$ $\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$ is defined to represent a noisy version of the depth map in the latent space at diffusion timestep $t$. The ground truch depth map $y^{*}$ is encoded into a latent representation $\mathbf{z_{y}}$ as supervision using a frozen VAE encoder $E$: $\mathbf{z_{y}}=E(y^{*})$. The diffusion latent after the whole denoising process $\mathbf{z_{0}}$ is decoded back to the image space using the frozen VAE decoder $D$: $\hat{y}=D(\mathbf{z}_{0})$.
@@ 4 Experiments
Training datasets. We adopt the training scheme of Marigold [30], and train our PriorDiffusion on two synthetic datasets that cover both indoor and outdoor scenes. The first dataset utilized is HyperSim [53], a photorealistic compilation of 461 indoor scenes. From HyperSim, approximately 54,000 samples is selected from 365 scenes, adhering to the official data split and excluding any incomplete samples. RGB images and their corresponding depth maps are resized to a resolution of 480 × 640 pixels. The second dataset is Virtual KITTI [3, 18], which comprises synthetic street scenes across five distinct settings, with varied conditions such as different weather patterns and camera perspectives. We specifically use the Virtual KITTI 2 [3] version. For training, we select four scenes, totaling approximately 20,000 samples.
@@
Implementation details. We follow the setup of Marigold [30]. Our method is implemented using PyTorch, using Stable Diffusion v2 [54] as the backbone model. During training, we use the DDPM noise scheduler [26] with 1000 diffusion steps. For inference, we use the DDIM scheduler [61] with 50 sampling steps. Our training process spans 30,000 iterations, with an effective batch size of 32 achieved through gradient accumulation over 16 steps (each with a per-step batch size of 2) to facilitate single-GPU execution. Optimization is performed using the Adam optimizer with a learning rate of $3\cdot 10^{-5}$. Additionally, random horizontal flipping is applied as a data augmentation. Training to convergence, which typically occurs after 20,000 iterations for our method, required approximately four days on an Nvidia RTX 3090 GPU.
@@ C.2. Ablation for prompts to generate text
This prompt is designed to elicit responses that include essential details, such as the positioning of objects, their interactions, and notable features that may impact depth estimation. Note that all training data we use are synthetic data, and by emphasizing “assuming it’s a real-world image,” we ensure that the descriptions align with the types of inputs and scenarios the model will encounter.